The director of a hospital wishes to estimate the mean number of people who are admitted to the emergency room during a 24-hour period. The director randomly selects 64 different 24-hour periods and determines the number of admissions for each. For this sample, sample mean is 19.5 and standard deviation is 25. Estimate the mean number of admissions per 24 hours period with 90% confidence interval.
Answer #1

Solution :- given that. x = 19.5 s = 25 n = 64 degree of freedom = n-1= 64-1 = 63 At go% confidence level the tis LET - 90 x=